OpenStudy (volpina):

Find the sum of the values satisfying the Mean Value Theorem given the function f(x)=(x^3/3)-2x^2+3x+4 on the interval 0≤x≤4 @3mar do you think you could help me?

OpenStudy (3mar):

What do you know about "the Mean Value Theorem"?

OpenStudy (volpina):

Not much... I know that the slope is equal to f(b)-f(a)/b-a

OpenStudy (3mar):

That is right .. and this is the complete view of it! http://www.sosmath.com/calculus/diff/der11/der11.html Can you plug in your values of f(0) and f(4) and find it out!?

OpenStudy (volpina):

f(0)=4 and f(4)= 16/3

OpenStudy (3mar):

and the ratio of \(\frac{f(b)-f(a)}{b-a}\)?

OpenStudy (3mar):

if that is your original function!! \[\LARGE f(x)=\frac{ 1 }{ 3 } x^3-2x^2+3x+4\]

OpenStudy (volpina):

1/3? and yes that's the original function

OpenStudy (3mar):

Very good! So you mean that: \[\Huge \frac{ f(4)-f(0) }{ 4-0 }=\frac{ \frac{ 16 }{ 3 } -4}{ 4-0 }=\color{Coral}{\frac{ 1 }{ 3 }}\] am I right?

OpenStudy (volpina):

yes

OpenStudy (3mar):

Congratulations! You hit the right target!

OpenStudy (volpina):

yay! thanks! could you help me with 2 more please?

OpenStudy (3mar):

Of course. With my pleasure!